y^4 • y ^7
x^3 • x^-5
j^a • j^b

Answer by algebrahouse.com(1659)

y^4(y^7)
= y^11 {add exponents when multiplying with same base}

x^3(x^-5)
= x^-2 {add exponents when multiplying with same base}
= 1/x² {made a positive exponent by moving to the denominator}

j^a(j^b)
= j^(a + b) {add exponents when multiplying with same base}

For more help from me, visit: www.algebrahouse.com